# List chromatic numbers and singular compactness

### v3

@@ 1. $(\aleph)$
Define $f\upharpoonright\bigcup_{\beta\in\alpha}A_{\beta}=\bigcup_{\beta\in\alpha}f\upharpoonright{A_{\beta}}$. Let $Y=A_{\alpha}-\bigcup_{\beta\in\alpha}A_{\beta}$, and notice that $Y$ can be empty. Since $Y\subseteq{A_{\alpha}}$ we see that ${\rm List}(Y)\leq{\rm List}(A_{\alpha})\leq\nu$. If $x\in{Y}$ then $|E^{x}\cap\bigcup_{\beta\in\alpha}A_{\beta}|<\nu$ since $\delta\leq{\rm cf}(\nu)$ and $\alpha\in\delta$. Thus one can define $F^{\prime}(x)=F(x)-(f^{\prime\prime}(E^{x}\cap\bigcup_{\beta\in\alpha}A_{\beta}))$, so $|F^{\prime}(x)|=\lambda$ and hence $F^{\prime}$ is a $\nu$-assignment for $Y$.
@@ 3. $(c)$
If ${\rm List}(H)\leq\lambda$ for every $H\leq G$ of size less than $\mu$ then ${\rm List}(G)\leq\lambda$. Hence if $\mu$ is strong limit then the statement holds for every $\lambda<\mu$.
@@
Let $(\mu_{i}:i\in\theta)$ be an increasing continuous sequence of cardinals such that $\theta,\lambda<\mu_{0}$ and $\mu=\bigcup_{i\in\theta}\mu_{i}$. For every $i\in\theta$ let $\nu_{i}=\mu_{i}^{\lambda}$, so $\nu_{i}<\mu$ for every $i\in\theta$ by virtue of $(b)$.
The following observation is central to the proof. If $U\subseteq V$ and $|U|=\mu_{i}$ then $|Y_{U}|<\nu_{i}$ where $Y_{U}=\{x\in V:|E^{x}\cap U|\geq\lambda\}$. Indeed, if $U$ forms a counterexample then one can apply Lemma 1 to the subgraph of $G$ induced by $U\cup Y_{U}$ and conclude that ${\rm List}(U\cup Y_{U})>\lambda$ contrary to the assumption that ${\rm List}(H)\leq\lambda$ whenever $H\leq G$ and $|H|<\mu$.
@@ 5. $(\varepsilon)$
    $G\in M^{\alpha}_{i}$ for every $i\in\theta,\alpha\in\lambda^{+}$.
@@
Pondering on the possibility that $2^{\lambda}<\mu$ is insufficient for singular compactness at $\mu$ (with $\lambda$ as a parameter) we conclude from the above theorem that in such a case there must be $\eta<\mu$ so that $\eta^{\lambda}\geq\mu$. The following claim shows that this cardinal $\eta$ has some interesting properties.
@@
Let $\eta$ be the first cardinal below $\mu$ for which $\eta^{\lambda}\geq\mu$. Such a cardinal exists by virtue of Theorem 4 and the assumptions of the claim. If $\theta<\mu$ and $\theta^{\lambda}\geq\eta$ then $\theta^{\lambda}=(\theta^{\lambda})^{\lambda}\geq\eta^{\lambda}\geq\mu$, and this is impossible since $\eta$ is the first cardinal which satisfies $\eta^{\lambda}\geq\mu$, hence $\theta<\eta\Rightarrow\theta^{\lambda}<\mu$. Thus necessarily ${\rm cf}(\eta)\leq\lambda$. Indeed, if ${\rm cf}(\eta)>\lambda$ then $\eta^{\lambda}=\bigcup_{\theta<\eta}\theta^{\lambda}=\eta<\mu$. The proof of the claim is accomplished. ∎
There are some interesting upshots which follow from this claim. Call $\mu$ a weird cardinal iff $\mu$ is singular and there is some $\lambda<\mu$ such that $2^{\lambda}<\mu$ and singular compactness for list chromatic numbers fails at $\mu$ with respect to $\lambda$. If $\lambda$ is specified then we shall say that $\mu$ is $\lambda$-weird. Call $\mu$ sane iff $\mu$ is not weird. Let us show that in some sense the collection of wierdos is small. Fix a cardinal $\kappa$ of uncountable cofinality. Suppose that $\mu<\kappa$ is $\lambda$-weird for some $\lambda<\mu$. From Claim 5 we see that there is some $\eta<\mu$ so that $\eta^{\lambda}\geq\mu$. If $\eta^{\lambda}<\kappa$ for every $\lambda$-wierd cardinal $\mu<\kappa$ then $\kappa$ will be called a sane cardinal.
@@
As a first step, we fix an infinite cardinal $\lambda$ and show that the set of $\lambda$-weird cardinals below $\kappa$ is not stationary. Call this set $S$, and assume towards contradiction that $S$ is stationary in $\kappa$. By cutting-off an initial segment we may assume that $2^{\lambda}<\min(S)$. Applying Claim 5 to each element of $S$, let $\eta^{\mu}<\mu$ be the pertinent cardinal, and since $\kappa$ is sane we see that $\eta_{\mu}^{\lambda}<\kappa$ for every $\eta\in S$.
The function $h(\mu)=\eta_{\mu}$ is regressive on $S$, hence there is a stationary subset $T\subseteq S$ and a fixed $\eta$ such that $\eta\in T\Rightarrow\eta_{\mu}=\eta$. Since $\kappa$ is sane we see that $\eta^{\lambda}<\kappa$. On the other hand, since $T$ is unbounded in $\kappa$ one can pick some $\mu\in T$ so that $\eta^{\lambda}<\mu$ and this is impossible since $\eta^{\lambda}=\eta_{\mu}^{\lambda}\geq\mu$.
@@ 1. $(\aleph)$
Usuba showed that ${\rm RP}({\rm List})$ has some consistency strength by proving that if ${\rm RP}({\rm List})$ holds then for every $\lambda\geq{\rm cf}(\lambda)>\omega$ either $2^{\lambda}>\lambda^{+}$ or every stationary subset of of $S^{\lambda^{+}}_{\omega}$ reflects. Consequently, either $\square_{\lambda}$ fails or SCH at $\lambda$ fails under ${\rm RP}({\rm List})$, see [Usu24, Proposition 4.6]. Another interesting fact is that ${\rm RP}({\rm List},2^{\omega})$ implies $2^{\omega}=\omega_{1}$. Based on these theorems, Usuba asked whether ${\rm RP}(\rm List)$ implies SCH. Our next goal is to prove that ${\rm GRP}({\rm List})$ implies SCH. In fact, ${\rm GRP}({\rm List})$ implies a bit more and this will be discussed after the proof. The proof itself is a generalization of Komjáth in [Kom13] who proved similar statements in the countable case.
@@
As a first step choose $\mathcal{A}=(A_{\alpha}:\alpha\in\mu)$ such that $|A_{\alpha}|=\mu$ for every $\alpha\in\mu$ and $\alpha_{0}\neq\alpha_{1}\Rightarrow A_{\alpha_{0}}\cap A_{\alpha_{1}}=\varnothing$. Now let $E=\prod\{A_{\alpha}:\alpha\in\mu\}$, so $|E|=\mu^{\mu}=2^{\mu}$. Define a $\mu$-assignment $F:A\cup B\rightarrow{\rm Ord}$ as follows. For every $\alpha\in\mu$ let $F(a_{\alpha})=A_{\alpha}$. Fix a bijection $h:B\rightarrow E$, say $h(b_{\beta})=g_{\beta}\in E$ for every $\beta\in 2^{\mu}$, and define $F(b_{\beta})=\{g_{\beta}(\alpha):\alpha\in\mu$. Notice that $|F(b_{\beta})|=\mu$ by the disjointness of the $A_{\alpha}$s.
Suppose that $f$ is a coloring defined on $A\cup B$ so that $f(v)\in F(v)$ for every $v\in A\cup B$. Since $f\upharpoonright{A}\in E$ there exists a unique $\beta\in 2^{\mu}$ such that $f\upharpoonright{A}=g_{\beta}$. It follows that $f(b_{\beta})\in F(b_{\beta})=\{g_{\beta}(\alpha):\alpha\in\mu\}=\{f(\alpha):\alpha\in\mu\}$, so for some $a_{\alpha}\in A$ we have $f(b_{\beta})=f(a_{\alpha})$. This fact shows that $f$ is not a good coloring since every element of $A$ is connected with every element of $B$, hence ${\rm List}(K_{\mu,2^{\mu}})>\mu$ as desired.

### v4

@@ 1. $(\aleph)$
Define $f\upharpoonright\bigcup_{\beta\in\alpha}A_{\beta}=\bigcup_{\beta\in\alpha}f\upharpoonright{A_{\beta}}$. Let $Y=A_{\alpha}-\bigcup_{\beta\in\alpha}A_{\beta}$, and notice that $Y$ can be empty. Since $Y\subseteq{A_{\alpha}}$ we see that ${\rm List}(Y)\leq{\rm List}(A_{\alpha})\leq\nu$. If $x\in{Y}$ then $|E^{x}\cap\bigcup_{\beta\in\alpha}A_{\beta}|<\nu$ since $\delta\leq{\rm cf}(\nu)$ and $\alpha\in\delta$. Thus one can define $F^{\prime}(x)=F(x)-(f^{\prime\prime}(E^{x}\cap\bigcup_{\beta\in\alpha}A_{\beta}))$, so $|F^{\prime}(x)|=\nu$ and hence $F^{\prime}$ is a $\nu$-assignment for $Y$.
@@ 3. $(c)$
If ${\rm List}(H)\leq\lambda$ for every $H\leq G$ of size less than $\mu$ then ${\rm List}(G)\leq\lambda$. Hence if $\mu$ is strong limit then the statement holds for every $\lambda\in{\rm Reg}\cap[\theta,\mu)$.
@@
Let $(\mu_{i}:i\in\theta)$ be an increasing continuous sequence of cardinals such that $\theta,\lambda<\mu_{0}$ and $\mu=\bigcup_{i\in\theta}\mu_{i}$. For every $i\in\theta$ let $\nu_{i}=\mu_{i}^{\lambda}$, so $\nu_{i}<\mu$ for every $i\in\theta$ by virtue of $(b)$. Notice that $\nu_{i}^{\lambda}=\nu_{i}$ for every $i\in\theta$.
The following observation is central to the proof. If $U\subseteq V$ and $|U|=\nu_{i}$ then $|Y_{U}|<\nu_{i}$ where $Y_{U}=\{x\in V:|E^{x}\cap U|\geq\lambda\}$. Indeed, if $U$ forms a counterexample then one can apply Lemma 1 to the subgraph of $G$ induced by $U\cup Y_{U}$ and conclude that ${\rm List}(U\cup Y_{U})>\lambda$ contrary to the assumption that ${\rm List}(H)\leq\lambda$ whenever $H\leq G$ and $|H|<\mu$.
@@ 5. $(\varepsilon)$
    $\mu,G\in M^{\alpha}_{i}$ for every $i\in\theta,\alpha\in\lambda^{+}$.
@@
Pondering upon the possibility that $2^{\lambda}<\mu$ is insufficient for singular compactness at $\mu$ (with $\lambda$ as a parameter) we conclude from the above theorem that in such a case there must be $\eta<\mu$ so that $\eta^{\lambda}\geq\mu$. The following claim shows that this cardinal $\eta$ has some interesting properties.
@@
Let $\eta$ be the first cardinal below $\mu$ for which $\eta^{\lambda}\geq\mu$. Such a cardinal exists by virtue of Theorem 4 and the assumptions of the claim. If $\theta<\eta$ and $\theta^{\lambda}\geq\eta$ then $\theta^{\lambda}=(\theta^{\lambda})^{\lambda}\geq\eta^{\lambda}\geq\mu$, and this is impossible since $\eta$ is the first cardinal which satisfies $\eta^{\lambda}\geq\mu$, hence $\theta<\eta\Rightarrow\theta^{\lambda}<\eta$. Thus necessarily ${\rm cf}(\eta)\leq\lambda$. Indeed, if ${\rm cf}(\eta)>\lambda$ then $\eta^{\lambda}=\bigcup_{\theta<\eta}\theta^{\lambda}=\eta<\mu$. The proof of the claim is accomplished. ∎
There are some interesting upshots which follow from this claim. Call $\mu$ a weird cardinal iff $\mu$ is singular and there is some $\lambda<\mu$ such that $2^{\lambda}\geq\mu$ and singular compactness for list chromatic numbers fails at $\mu$ with respect to $\lambda$. If $\lambda$ is specified then we shall say that $\mu$ is $\lambda$-weird. Let us show that in some sense the collection of wierdos is small. Fix a cardinal $\kappa$ of uncountable cofinality. Suppose that $\mu<\kappa$ is $\lambda$-weird for some $\lambda<\mu$. From Claim 5 we see that there is some $\eta<\mu$ so that $\eta^{\lambda}\geq\mu$. If $\eta^{\lambda}<\kappa$ for every $\lambda$-wierd cardinal $\mu<\kappa$ then $\kappa$ will be called a sane cardinal.
@@
As a first step, we fix an infinite cardinal $\lambda$ and show that the set of $\lambda$-weird cardinals below $\kappa$ is not stationary. Call this set $S$, and assume towards contradiction that $S$ is stationary in $\kappa$. By cutting-off an initial segment we may assume that $2^{\lambda}<\min(S)$. Applying Claim 5 to each element of $S$, let $\eta_{\mu}<\mu$ be the pertinent cardinal, and since $\kappa$ is sane we see that $\eta_{\mu}^{\lambda}<\kappa$ for every $\mu\in S$.
The function $h(\mu)=\eta_{\mu}$ is regressive on $S$, hence there is a stationary subset $T\subseteq S$ and a fixed $\eta$ such that $\mu\in T\Rightarrow\eta_{\mu}=\eta$. Since $\kappa$ is sane we see that $\eta^{\lambda}<\kappa$. On the other hand, since $T$ is unbounded in $\kappa$ one can pick some $\mu\in T$ so that $\eta^{\lambda}<\mu$ and this is impossible since $\eta^{\lambda}=\eta_{\mu}^{\lambda}\geq\mu$.
@@ 1. $(\aleph)$
Usuba showed that ${\rm RP}({\rm List})$ has some consistency strength by proving that if ${\rm RP}({\rm List})$ holds then for every $\lambda\geq{\rm cf}(\lambda)>\omega$ either $2^{\lambda}>\lambda^{+}$ or every stationary subset of $S^{\lambda^{+}}_{\omega}$ reflects. Consequently, either $\square_{\lambda}$ fails or SCH at $\lambda$ fails under ${\rm RP}({\rm List})$, see [Usu24, Proposition 4.6]. Another interesting fact is that ${\rm RP}({\rm List},2^{\omega})$ implies $2^{\omega}=\omega_{1}$. Based on these theorems, Usuba asked whether ${\rm RP}(\rm List)$ implies SCH. Our next goal is to prove that ${\rm GRP}({\rm List})$ implies SCH. In fact, ${\rm GRP}({\rm List})$ implies a bit more and this will be discussed after the proof. The proof itself is a generalization of Komjáth in [Kom13] who proved similar statements in the countable case.
@@
As a first step choose $\mathcal{A}=(A_{\alpha}:\alpha\in\mu)$ such that $|A_{\alpha}|=\mu$ for every $\alpha\in\mu$ and $\alpha_{0}\neq\alpha_{1}\Rightarrow A_{\alpha_{0}}\cap A_{\alpha_{1}}=\varnothing$. Now let $E=\prod\{A_{\alpha}:\alpha\in\mu\}$, so $|E|=\mu^{\mu}=2^{\mu}$. Define a $\mu$-assignment $F:A\cup B\rightarrow[{\rm Ord}]^{\mu}$ as follows. For every $\alpha\in\mu$ let $F(a_{\alpha})=A_{\alpha}$. Fix a bijection $h:B\rightarrow E$, say $h(b_{\beta})=g_{\beta}\in E$ for every $\beta\in 2^{\mu}$, and define $F(b_{\beta})=\{g_{\beta}(\alpha):\alpha\in\mu\}$. Notice that $|F(b_{\beta})|=\mu$ by the disjointness of the $A_{\alpha}$s.
Suppose that $f$ is a coloring defined on $A\cup B$ so that $f(v)\in F(v)$ for every $v\in A\cup B$. Since $f\upharpoonright{A}\in E$ there exists a unique $\beta\in 2^{\mu}$ such that $f\upharpoonright{A}=g_{\beta}$. It follows that $f(b_{\beta})\in F(b_{\beta})=\{g_{\beta}(\alpha):\alpha\in\mu\}=\{f(a_{\alpha}):\alpha\in\mu\}$, so for some $a_{\alpha}\in A$ we have $f(b_{\beta})=f(a_{\alpha})$. This fact shows that $f$ is not a good coloring since every element of $A$ is connected with every element of $B$, hence ${\rm List}(K_{\mu,2^{\mu}})>\mu$ as desired.